# When Large Language Models Meet UAVs: How Far Are We?

### v1

@@ 7.1. UAV-LLM Integration
A growing body of research has investigated concrete systems and technical frameworks for integrating LLMs with UAVs. Chen et al.(Chen et al., 2023) developed TypeFly, the first real-world UAV platform integrated with an LLM to support natural language-driven flight control. Their system demonstrated the feasibility of mapping natural language to motion commands but also revealed limitations in handling semantic ambiguity and adapting to dynamic environments. Phadke et al.(Phadke et al., 2024) proposed a modular UAV control framework based on LLMs in simulation, coordinating navigation, obstacle avoidance, and target recognition modules. While the modular design improved stability, the absence of a unified semantic memory layer restricted decision efficiency. Zhao and Lin (Zhao et al., 2023b) introduced the concept of a Universal Aerial Agent with the “Prompt–as–Policy” approach, enabling flexible encoding of task intents, though their work remains at a conceptual stage without large-scale validation. More recently, Koubaa (Koubaa et al., 2025) introduced UAV–GPT, which translates high-level semantic instructions into executable UAV tasks, while Ren et al. (Ren et al., 2024) integrated foundation models with UAV perception for semantic annotation and visual reasoning.
@@
These technical explorations highlight the potential of LLMs to enhance UAV autonomy, perception, and interaction, yet most remain isolated prototypes or proof-of-concept systems. Compared to these works, our study provides a more comprehensive and integrative perspective: we categorize task roles across the UAV workflow, empirically compare academic versus industrial tendencies, and identify bottlenecks from practitioner data. This broader view helps to contextualize existing prototypes within a systematic landscape and points out where research and practice can be better aligned.
7.2. Surveys on UAVs and Machine Learning
@@
A number of surveys have summarized the use of machine learning techniques in UAVs. For example, Osco et al.(Osco et al., 2021) reviewed the role of deep learning in UAV remote sensing, covering tasks such as image classification, object recognition, and environmental monitoring. Tang et al.(Tang et al., 2023) provided a survey on object detection for UAVs based on deep learning, emphasizing algorithmic advances and benchmark datasets. AlMahamid and Grolinger (AlMahamid and Grolinger, 2022) systematically reviewed reinforcement learning approaches for autonomous UAV navigation, highlighting progress and challenges in dynamic environments. Bai et al. (Bai et al., 2023) focused on reinforcement learning–based methods for building autonomous multi-UAV wireless networks, summarizing techniques for coordination and communication. These surveys clearly show the important role of machine learning techniques in UAV research, but none of them address the emerging opportunities and challenges brought by large language models. In other words, although UAV–DL/ML integration has been extensively reviewed, the UAV–LLM integration landscape remains largely unexplored.
@@
There has been one recent attempt to discuss UAV-LLM integration. Javaid et al. (Javaid et al., 2024) reviewed possible applications of LLMs in UAVs, focusing on natural language command control, reasoning-assisted decision making, and human-computer interaction. However, their work mainly focuses on conceptual and descriptive descriptions, without connecting academic, industry, and developers’ perspectives. In contrast, our study constructs the first taxonomy of UAV-LLM tasks, analyzes 997 academic papers and 1,509 open-source projects, and further validates our findings with survey responses from 52 developers. This large-scale effort not only summarizes existing research but also quantifies the gap between academia and industry, offering practical guidance for real-world deployment.
[…]
8. Conclusion
@@
In this paper, we conduct an empirical study on the integration of LLMs and UAVs to examine the current state of applications in academia and industry. First, by analyzing 74 papers and 56 open-source projects on GitHub, we summarize nine categories of LLM-supported tasks across four UAV workflows. Second, we quantify the distribution of these tasks in academic and industrial settings and perform a comparative analysis. The results reveal gaps between academia and industry in research focus, technical requirements, and integration modes of LLMs and UAVs. We then administer an online survey and obtain 52 valid responses. We find that 59.6% of practitioners have not yet attempted to use LLMs in real-world UAV projects. Their feedback indicates that current LLMs fall short of meeting the requirements of UAVs for accuracy, real-time performance, and reliability, and that the lack of technological maturity and practical feasibility in academic research further increases the complexity of engineering deployment. Finally, based on our findings, we highlight future challenges for UAV-LLM integration and offer recommendations for researchers and practitioners.

### v2

@@ 7.1 UAV-LLM Integration
A growing body of research has investigated concrete systems and technical frameworks for integrating LLMs with UAVs. Chen et al.[chen2023typefly] developed TypeFly, the first real-world UAV platform integrated with an LLM to support natural language-driven flight control. Their system demonstrated the feasibility of mapping natural language to motion commands but also revealed limitations in handling semantic ambiguity and adapting to dynamic environments. Phadke et al.[phadke2024integrating] proposed a modular UAV control framework based on LLMs in simulation, coordinating navigation, obstacle avoidance, and target recognition modules. While the modular design improved stability, the absence of a unified semantic memory layer restricted decision efficiency. Zhao and Lin [zhao2025general] introduced the concept of a Universal Aerial Agent with the “Prompt–as–Policy” approach, enabling flexible encoding of task intents, though their work remains at a conceptual stage without large-scale validation. More recently, Koubaa [koubaa2025next] introduced UAV–GPT, which translates high-level semantic instructions into executable UAV tasks, while Ren et al. [ren2024grounded] integrated foundation models with UAV perception for semantic annotation and visual reasoning.
In addition to these LLM-focused studies, several works have examined UAV software quality and system reliability, which are becoming increasingly important as UAVs adopt more intelligent control strategies. Di Sorbo et al.[di2023automated]analyzed safety concerns reported in UAV software platforms and showed that many issues stem from fragile system design. Shar et al.[shar2022dronlomaly]introduced DronLomaly, a method for detecting abnormal drone behaviors at runtime through deep log analysis, demonstrating the value of continuous monitoring. Wang et al.[wang2025routhsearch]proposed RouthSearch to infer PID parameter specifications for flight-control programs, helping improve the robustness of low-level control. Furthermore, Wang et al.[wang2024exploratory, wang2021exploratory]conducted exploratory studies on UAV log anomalies and autopilot software bugs, offering empirical insights into common reliability problems in existing UAV systems.
@@
These technical explorations highlight the potential of LLMs to enhance the autonomy, perception, and interaction of UAVs. However, most existing systems are still early prototypes without large-scale validation. The software-engineering studies discussed above also indicate that safety, anomaly detection, and reliable control are necessary foundations for practical UAV-LLM integration. Compared to these works, our study provides a more comprehensive perspective: we categorize tasks across the UAV workflow, compare differences between academic and industrial practices, and identify key challenges reported by developers. This broader view helps to contextualize existing prototypes within a systematic landscape and points out where research and practice can be better aligned.
7.2 Empirical Studies on UAVs
@@
A number of empirical studies have summarized the use of machine learning techniques in UAVs. For example, Osco et al.[osco2021review] reviewed the role of deep learning in UAV remote sensing, covering tasks such as image classification, object recognition, and environmental monitoring. Tang et al.[tang2023survey] provided a survey on object detection for UAVs based on deep learning, emphasizing algorithmic advances and benchmark datasets. AlMahamid and Grolinger [almahamid2022autonomous] systematically reviewed reinforcement learning approaches for autonomous UAV navigation, highlighting progress and challenges in dynamic environments. Bai et al. [bai2023toward] focused on reinforcement learning–based methods for building autonomous multi-UAV wireless networks, summarizing techniques for coordination and communication. These surveys clearly show the important role of machine learning techniques in UAV research, but none of them address the emerging opportunities and challenges brought by large language models. In other words, although UAV–DL/ML integration has been extensively reviewed, the UAV–LLM integration landscape remains largely unexplored.
@@
There has been one recent attempt to discuss UAV-LLM integration. Javaid et al. [javaid2024large] reviewed possible applications of LLMs in UAVs, focusing on natural language command control, reasoning-assisted decision making, and human-computer interaction. However, their work mainly focuses on conceptual and descriptive descriptions, without connecting academic, industry, and developers’ perspectives. In contrast, our study constructs the first taxonomy of UAV-LLM tasks, analyzes 997 academic papers and 1,509 open-source projects, and further validates our findings with survey responses from 52 developers. This large-scale effort not only summarizes existing research but also quantifies the gap between academia and industry, offering practical guidance for real-world deployment.
8 Conclusion
@@
In this paper, we conduct an empirical study of LLM applications in UAVs to examine the current state of research and development. First, by analyzing 74 papers and 56 open-source GitHub projects, we summarize nine categories of LLM-supported tasks across four UAV workflows. Second, we quantify the distribution of these tasks in researchers’ and developers’ practices and perform a comparative analysis. The results reveal gaps between researchers and developers in research focus, technical requirements, and integration modes of LLMs and UAVs. We then administer an online survey and obtain 52 valid responses. We find that 59.6% of developers have not yet attempted to use LLMs in real-world UAV projects. Their feedback indicates that current LLMs fall short of meeting the requirements of UAVs for accuracy, real-time performance, and reliability, and that the lack of technical maturity and practical feasibility in academic research further increases the complexity of engineering deployment. Finally, based on our findings, we highlight future challenges for the application of LLMs in UAVs and offer recommendations for researchers and developers.